# Distributed Algorithms for Potential Problems

### v3

@@
Abstract. In this work we present a fast distributed algorithm for local potential problems: these are graph problems where the task is to find a locally optimal solution where no node can unilaterally improve the utility in its local neighborhood by changing its own label. A simple example of such a problem is the task of finding a locally optimal cut, i.e., a cut where for each node at least half of its incident edges are cut edges.
The distributed round complexity of locally optimal cut has been wide open; the problem is known to require $\Omega(\operatorname{log}n)$ rounds in the deterministic LOCAL model and $\Omega(\operatorname{log}\operatorname{log}n)$ rounds in the randomized LOCAL model, but the only known upper bound is the trivial brute-force solution of $O(n)$ rounds. Locally optimal cut in bounded-degree graphs is perhaps the simplest example of a locally checkable labeling problem for which there is still such a large gap between current upper and lower bounds.
We show that in bounded-degree graphs, all local potential problems, including locally optimal cut, can be solved in $\operatorname{log}^{O(1)}n$ rounds, both in the deterministic and randomized LOCAL models. In particular, the deterministic round complexity of the locally optimal cut problem is now settled to $\operatorname{log}^{\Theta(1)}n$.
@@ 1 Introduction
In this work, we present a distributed algorithm that finds a locally optimal cut in any bounded-degree graph in a polylogarithmic number of rounds (see Section 1.1); this is the first nontrivial distributed algorithm for this problem. Moreover, our algorithm generalizes to all distributed graph problems that are defined with a local potential function (see Section 1.2).
@@ 1.1 Locally optimal cuts
While it is trivial to find a locally optimal cut with a centralized algorithm, in this work we are interested in distributed algorithms for finding a locally optimal cut. We work in the usual LOCAL model [Lin92, Pel00] of distributed computing: Each node $v\in V$ represents a computer and each edge $\{u,v\}$ represents a communication link between two computers. Initially each node is only aware of its own unique identifier, its degree, and the total number of nodes $n$. Computation proceeds in synchronous rounds; in each round each node can send a message to each neighbor, receive a message from each neighbor, and update its own state. Eventually each node $v$ has to stop and output its own part of the solution, here $\ell(v)$. The running time of the algorithm is $T(n)$ if in any $n$-node graph all nodes stop after at most $T(n)$ rounds.
@@
It is not hard to turn the trivial centralized algorithm into a valid distributed algorithm for solving the locally optimal cut problem. For example, identify all unhappy nodes, then find a maximal independent set of unhappy nodes, and flip all of them in parallel. Unfortunately, this is not going to be an efficient algorithm. Consider the following graph and a cut $\ell$ indicated with the colors white and black; here the node marked with an arrow is the only unhappy node, as only $1$ of its $3$ incident edges are cut edges:
@@
Flipping that we again move unhappiness around:
Clearly this will take $\Omega(n)$ rounds until all unhappiness is resolved (even if we put aside the time needed to find a maximal independent set), so it is no better than the trivial brute-force solution that solves any graph problem in $O(n)$ rounds. Can we do much better than this?
@@
We will study this question in bounded-degree graphs, i.e., we assume that there is some fixed constant $\Delta$ and our goal is to solve locally optimal cut in any graph with maximum degree at most $\Delta$. We note that e.g. in the above apparently-hard instance we had $\Delta=3$, so small values of $\Delta$ do not make the problem trivial.
For any fixed $\Delta$, the locally optimal cut problem is an example of a locally checkable labeling problem (LCL). LCL problems were first defined by [NS95], and there is a large body of work that has developed a complexity theory of LCL problems in the LOCAL model [CV86, Nao91, Lin92, NS95, BFH+16, FG17, GHK18, BHK+18, CP19, CKP19, RG20, BBO+20, BBO+21, DDL+23, Cha20, BBO+21, GRB22, Suo20].
Most natural LCL problems are by now very well-understood, especially for the deterministic LOCAL model. Indeed, for some restricted families of LCLs and restricted families of graphs, there is now a complete classification of the round complexity of all such problems, see e.g. [BBE+20, CSS23, BHK+17, BBC+22, BBC+19]. For example, there is a broad class of symmetry-breaking problems, all of which have round complexity $\Theta(\operatorname{log}^{*}n)$ [CV86, GPS88, Lin92, Nao91], and many other problems [GHK+20] have turned out to be asymptotically as hard as the sinkless orientation problem [BFH+16], which has round complexity $\Theta(\operatorname{log}n)$ in the deterministic LOCAL model [BFH+16, CKP19, GS17].
Locally optimal cut is perhaps the most elementary LCL problem whose distributed complexity is currently wide open. We do not even know if it admits any nontrivial algorithm with round complexity $o(n)$, not even for the smallest nontrivial degree bound $\Delta=3$. (The case of $\Delta=2$ is fully understood, as it is in essence equivalent to the weak-$2$-coloring problem, which can be solved in $O(\operatorname{log}^{*}n)$ rounds, and this is also tight, see e.g. [BHO+19].)
@@
What makes locally optimal cuts particularly intriguing is that many commonly-used lower bound techniques fail. In particular, one cannot use existential graph-theoretic arguments (in the spirit of e.g. the lower bound for coloring trees in [Lin92] or the lower bound for approximate coloring in [CDG+24]), since a locally optimal cut always exists in any graph.
It is, however, possible to construct a nontrivial reduction that connects locally optimal cuts with the sinkless orientation problem [BHL+19], and this gives the following lower bounds: we need at least $\Omega(\operatorname{log}n)$ rounds in the deterministic LOCAL model, and at least $\Omega(\operatorname{log}\operatorname{log}n)$ rounds in the randomized LOCAL model.
This is essentially everything that is known about locally optimal cuts before this work. There are no upper bounds better than the trivial $O(n)$. Furthermore, the lower bounds do not hold in stronger models such as quantum-LOCAL [GKM09, AF14, ACD+25, BBC+25]—there we do not even know if the true complexity of the problem is closer to $O(1)$ or $\Omega(n)$.
We note that locally optimal cuts have been studied e.g. in the context of descriptive combinatorics also under the names of unfriendly colorings and unfriendly partitions [SM12, Con14, CMU20, CT20].
@@
We present the first nontrivial upper bound for locally optimal cuts. We show that the problem can be solved in $\operatorname{log}^{O(1)}n$ rounds, both in the deterministic and randomized LOCAL models. In particular, the complexity of locally optimal cut in the deterministic LOCAL model is now known to be $\operatorname{log}^{\Theta(1)}n$ rounds.
@@
The main open question is related to models stronger than deterministic LOCAL. In the randomized LOCAL model, there is still an exponential gap between the upper and lower bounds, and in the quantum-LOCAL model we are lacking any nontrivial lower bounds.
@@ 1.2 Local potential problems
At first the definition of locally optimal problems may feel very restrictive, as we only consider solutions where no single node can change its own label to improve the potential. In many problems we would like to make more complicated local changes (e.g. consider the maximum independent set problem, and the task of finding a locally optimal solution where there is no place where removing one node would let us add two nodes). However, in the bounded-degree setting we can easily encode such problems in our formalism. To see this, assume we have a graph problem with $k$ labels. Then the local label of a node $v$ can be a vector of $k^{\Delta+1}$ labels $\ell_{0},\dotsc,\ell_{\Delta}$, where $\ell_{0}$ is the “base” label of node $v$, and $\ell_{i}$ indicates that node $v$ wants to add $\ell_{i}$ modulo $k$ to the label of its $i$th neighbor. This way changing the physical label of one node allows us to change the logical labels of all nodes in its radius-$1$ neighborhood. This scheme of course generalizes also to any constant radius. Hence without loss of generality it suffices to consider solutions where no single node can unilaterally improve the total utility.
@@
In this work we design an algorithm that solves any locally optimal problem in $\operatorname{log}^{O(1)}n$ rounds in bounded-degree graphs in the deterministic LOCAL model. As a corollary, we can also solve any local potential problem in the same time. Our result for locally optimal cuts now follows as a special case of this.
@@
Conversely, super-polylogarithmic lower bounds on the distributed round complexity now directly imply that a problem cannot be represented with the help of a local potential. To give a simple example, stable matchings require a linear number of rounds, even in bounded-degree bipartite graphs [FKP+10], so this problem cannot be represented with local potential functions. Results of this form may be of interest also in a much broader context outside distributed computing.
@@
Key examples of local potential problems include defective coloring problems, see e.g. [BE13, Chapter 6], [CPS17, FG17]. Assume that we have $c$ colors and we want to have defect at most $d$, that is, each node can have at most $d$ neighbors with the same color. If $c(d+1)>\Delta$, this is a local potential problem. Here the potential function is the number of properly colored edges. If a node is unhappy (it has more than $d$ neighbors with the same color), it can find a color that makes it happy. Our algorithm can be used to solve this problem in for any constants $c$, $d$ and $\Delta$ with $c(d+1)>\Delta$ in polylogarithmic number of rounds.
@@
A key open question is understanding the distributed complexity landscape of the family of local potential problems. Our work together shows that there are local potential problems with deterministic round complexity $\operatorname{log}^{\Theta(1)}n$, and by prior work there are also local potential problems with round complexity $\Theta(\operatorname{log}^{*}n)$ and $O(1)$. However, one question that is currently wide open is what happens when we step outside the deterministic LOCAL model. For example, does randomness ever help with any of these problems? The only lower bound we have is the exponentially lower $\Omega(\operatorname{log}\operatorname{log}n)$ for locally optimal cuts, is this the correct complexity? If yes do all local potential problems admit $\operatorname{poly}(\operatorname{log}\operatorname{log}n)$ randomized algorithms? Do any of these problems admit a distributed quantum advantage?
Also questions related to decidability are wide open. Fore example, given the specification of a local potential problem, is it decidable whether it can be solved in $O(\operatorname{log}^{*}n)$ rounds or whether it requires $\Omega(\operatorname{log}n)$ rounds?
Finally, our work only considered the case of bounded-degree graphs. Can we solve local potential problems in polylogarithmic time also in general graphs?
@@ 1.3 New techniques and key ideas
Intuitively, the key challenge in our example was that the solution was very fragile—all nodes were barely happy, and a minor perturbation in their local neighborhood may result them to flip, which leads to long cascades that we must avoid.
@@
Let us now try to identify a key combinatorial property of locally optimal cuts (and more generally, local potential problems). Consider first a sequential algorithm that eliminates all unhappy nodes one by one, let $X$ be the set of all nodes that were flipped in this process at least once. Note that in general, $X$ might span the entire input graph.
@@
Our algorithm uses the MPX clustering algorithm [MPX13, EN22] to divide the nodes into polylogarithmic-radius clusters with small boundaries. Each cluster can make nodes far from boundaries safe; we can simply compute for each cluster the result of repeatedly finding for each internal node an improving set $A$, and flipping $A$ until all internal nodes are safe.
@@
Our algorithm is randomized, but it can be derandomized (with polylogarithmic cost) by using the standard techniques [GHK18, RG20].

### v4

@@
Abstract. In this work, we present a fast distributed algorithm for local potential problems: these are graph problems where the task is to find a locally optimal solution where no node can unilaterally improve the utility in its local neighborhood by changing its own label. A simple example of such a problem is the task of finding a locally optimal cut, i.e., a cut where for each node at least half of its incident edges are cut edges.
The distributed round complexity of the locally optimal cut problem has been wide open; the problem is known to require $\Omega(\operatorname{log}n)$ rounds in the deterministic LOCAL model and $\Omega(\operatorname{log}\operatorname{log}n)$ rounds in the randomized LOCAL model, but the only known upper bound is the trivial brute-force solution of $O(n)$ rounds. Locally optimal cut in constant-degree graphs is perhaps the simplest example of a locally checkable labeling problem for which there is still such a large gap between current upper and lower bounds.
We show that in constant-degree graphs, all local potential problems, including locally optimal cut, can be solved in $\operatorname{log}^{O(1)}n$ rounds, both in the deterministic and randomized LOCAL models. In particular, the deterministic round complexity of the locally optimal cut problem is now settled to $\operatorname{log}^{\Theta(1)}n$.
Our algorithms also apply to the general case of graphs of maximum degree $\Delta$. For the special case of locally optimal cut, we obtain a randomized algorithm that runs in $O(\Delta^{2}\operatorname{log}^{6}n)$ rounds, which can be derandomized at polylogarithmic cost with standard techniques. Furthermore, we show that a dependence in $\Delta$ is necessary: we prove a lower bound of $\Omega(\min\{\Delta,\sqrt{n}\})$ rounds, even in the quantum-LOCAL model; in particular, there is no polylogarithmic-round algorithm for the general case.
@@ 1 Introduction
In this work, we present a distributed algorithm that finds a locally optimal cut in any constant-degree graph in a polylogarithmic number of rounds (see Section 1.1); this is the first nontrivial distributed algorithm for this problem. Moreover, our algorithm generalizes to all distributed graph problems that are defined with a local potential function (see Section 1.2).
@@ 1.1 Locally optimal cuts
While it is trivial to find a locally optimal cut with a centralized algorithm, in this work we are interested in distributed algorithms for finding a locally optimal cut. We work in the usual LOCAL model [39, 43] of distributed computing: Each node $v\in V$ represents a computer and each edge $\{u,v\}$ represents a communication link between two computers. Initially each node is only aware of its own unique identifier, its degree, and the total number of nodes $n$. Computation proceeds in synchronous rounds; in each round each node can send a message to each neighbor, receive a message from each neighbor, and update its own state. Eventually each node $v$ has to stop and output its own part of the solution, here $\ell(v)$. The running time of the algorithm is $T(n)$ if in any $n$-node graph all nodes stop after at most $T(n)$ rounds.
@@
It is not hard to turn the trivial centralized algorithm into a valid distributed algorithm for solving the locally optimal cut problem. For example, identify all unhappy nodes, then find a maximal independent set of unhappy nodes, and flip all of them in parallel. Unfortunately, this is not going to be an efficient algorithm. Consider the following graph and a cut $\ell$ indicated with the colors white and black; here the node marked with an arrow is the only unhappy node, as only $1$ of its $3$ incident edges is a cut edge:
@@
Flipping that node, we again move unhappiness around:
Clearly, this will take $\Omega(n)$ rounds until all unhappiness is resolved (even if we put aside the time needed to find a maximal independent set), so it is no better than the trivial brute-force solution that solves any graph problem in $O(n)$ rounds. Can we do much better than this?
@@
The problem of finding a locally optimal cut is non-trivial even when restricting to bounded-degree graphs, i.e., when there is some fixed constant $\Delta$ and our goal is to solve locally optimal cut in any graph with maximum degree at most $\Delta$. We note that e.g. in the above apparently-hard instance we had $\Delta=3$, so small values of $\Delta$ do not make the problem trivial.
For any fixed $\Delta$, the locally optimal cut problem is an example of a locally checkable labeling problem (LCL). LCL problems were first defined by [41], and there is a large body of work that has developed a complexity theory of LCL problems in the LOCAL model [23, 42, 39, 41, 15, 29, 33, 11, 18, 17, 44, 7, 8, 27, 20, 8, 38, 46].
Most natural LCL problems are by now very well-understood, especially for the deterministic LOCAL model. Indeed, for some restricted families of LCLs and restricted families of graphs, there is now a complete classification of the round complexity of all such problems, see e.g. [6, 19, 16, 4, 3]. For example, there is a broad class of symmetry-breaking problems, all of which have round complexity $\Theta(\operatorname{log}^{*}n)$ [23, 37, 39, 42], and many other problems [34] have turned out to be asymptotically as hard as the sinkless orientation problem [15], which has round complexity $\Theta(\operatorname{log}n)$ in the deterministic LOCAL model [15, 17, 36].
Locally optimal cut is perhaps the most elementary LCL problem whose distributed complexity is currently wide open. We do not even know if it admits any nontrivial algorithm with round complexity $o(n)$, not even for the smallest nontrivial degree bound $\Delta=3$. (The case of $\Delta=2$ is fully understood, as it is in essence equivalent to the weak-$2$-coloring problem, which can be solved in $O(\operatorname{log}^{*}n)$ rounds, and this is also tight, see e.g. [13].)
@@
What makes locally optimal cuts particularly intriguing is that many commonly-used lower bound techniques fail. In particular, one cannot use existential graph-theoretic arguments (in the spirit of e.g. the lower bound for coloring trees in [39] or the lower bound for approximate coloring in [22]), since a locally optimal cut always exists in any graph.
It is, however, possible to construct a nontrivial reduction that connects locally optimal cuts with the sinkless orientation problem [12], and this gives the following lower bounds: we need at least $\Omega(\operatorname{log}n)$ rounds in the deterministic LOCAL model, and at least $\Omega(\operatorname{log}\operatorname{log}n)$ rounds in the randomized LOCAL model.
This is essentially everything that is known about locally optimal cuts before this work. There are no upper bounds better than the trivial $O(n)$. Furthermore, the lower bounds do not hold in stronger models such as quantum-LOCAL [31, 2, 1, 5, 9, 10]—there we do not even know if the true complexity of the problem is closer to $O(1)$ or $\Omega(n)$.
We note that locally optimal cuts have been studied e.g. in the context of descriptive combinatorics also under the names of unfriendly colorings and unfriendly partitions [45, 26, 24, 25].
@@
We present the first nontrivial upper bound for locally optimal cuts. We show that, in bounded-degree graphs, the problem can be solved in $\operatorname{log}^{O(1)}n$ rounds, both in the deterministic and randomized LOCAL models. In particular, the complexity of locally optimal cut in the deterministic LOCAL model is now known to be $\operatorname{log}^{\Theta(1)}n$ rounds.
In the case where the maximum degree $\Delta$ of the graph is not constant, our algorithm requires $O(\Delta^{2}\operatorname{log}^{6}n)$ randomized rounds (see Corollary 3.8 for the main result), which can be derandomized at polylogarithmic cost by using standard techniques [33, 44, 32], obtaining a deterministic runtime of $O(\Delta^{2}\operatorname{log}^{8}(n)\operatorname{poly}(\operatorname{log}\operatorname{log}n))$ (Corollary 3.13). Furthermore, we show that $\Omega(\min\{\Delta,\sqrt{n}\})$ rounds are necessary, even for randomized and quantum algorithms; in particular, there is no polylogarithmic-round algorithm for the general case (see Theorem 5.2 for the result).
@@
The complexity of the problem is not completely settled yet. For bounded-degree graphs, the situation is the following: In the deterministic LOCAL model there is now a polynomial (in $\operatorname{log}n$) gap between the upper and lower bounds. In the randomized LOCAL model, there is still an exponential gap between the upper and lower bounds, and in the quantum-LOCAL model we are lacking any nontrivial lower bounds. For unbounded-degree graphs, there is still a polynomial gap between the upper and lower bounds, even in the quantum-LOCAL model.
@@ 1.2 Local potential problems
At first the definition of locally optimal problems may feel very restrictive, as we only consider solutions where no single node can change its own label to improve the potential. In many problems we would like to make more complicated local changes (e.g. consider the maximum independent set problem, and the task of finding a locally optimal solution where there is no place where removing one node would let us add two nodes). However, in the bounded-degree setting we can easily encode such problems in our formalism. To see this, assume we have a graph problem with $k$ labels. Then the local label of a node $v$ can be a vector of $k^{\Delta+1}$ labels $\ell_{0},\dotsc,\ell_{\Delta}$, where $\ell_{0}$ is the “base” label of node $v$, and $\ell_{i}$ indicates that node $v$ wants to add $\ell_{i}$ modulo $k$ to the label of its $i$th neighbor. This way changing the physical label of one node allows us to change the logical labels of all nodes in its radius-$1$ neighborhood. This scheme of course generalizes also to any constant radius. Hence, without loss of generality, it suffices to consider solutions where no single node can unilaterally improve the total utility.
@@
In this work we design an algorithm that solves any locally optimal problem in $\operatorname{log}^{O(1)}n$ rounds in bounded-degree graphs in the deterministic LOCAL model. As a corollary, we can also solve any local potential problem in the same time. We also analyze the dependence on the maximum degree $\Delta$, so that the results are directly applicable beyond the bounded-degree settings. In general, the randomized complexity of our algorithm is $\Delta^{O(1)}(\Lambda/\lambda)^{2}\operatorname{log}^{6}(\Lambda n/\lambda)$, where $\Lambda$ is the maximum improvement associated to the problem, and $\lambda$ is the minimum improvement associated to the problem: $\Lambda$ and $\lambda$ are defined by the sole description of the problem and the value of $\Delta$ (see Sections 2 and 3 for formal definitions, and Theorem 3.7 for the main result). Derandomizing the algorithm comes at a polylogarithmic cost, so the deterministic complexity is $\Delta^{O(1)}(\Lambda/\lambda)^{2}\operatorname{log}^{6}(\Lambda n/\lambda)\cdot\tilde{O}(\operatorname{log}^{2}(n))$ (Corollary 3.12). In particular, for any constant $\Delta$, it holds that $\Lambda=O(1)$ and $\lambda=O(1)$, and hence we have an $O(\operatorname{log}^{6}n)$-round randomized algorithm and an $\tilde{O}(\operatorname{log}^{8}n)$-round deterministic algorithm. Our result for locally optimal cuts now follows as a special case of this.
@@
Conversely, super-polylogarithmic lower bounds on the distributed round complexity now directly imply that an LCL problem cannot be represented with the help of a local potential. To give a simple example, stable matchings require a linear number of rounds, even in bounded-degree bipartite graphs [30], so this problem cannot be represented with local potential functions. Results of this form may be of interest also in a much broader context outside distributed computing.
@@
Key examples of local potential problems include defective coloring problems, see e.g. [14, Chapter 6], [21, 29]. Assume that we have $c$ colors and we want to have defect at most $d$, that is, each node can have at most $d$ neighbors with the same color. If $c(d+1)>\Delta$, this is a local potential problem. Here the potential function is the number of properly colored edges. If a node is unhappy (it has more than $d$ neighbors with the same color), it can find a color that makes it happy. Our algorithm can be used to solve this problem for any constants $c$, $d$ and $\Delta$ with $c(d+1)>\Delta$ in a polylogarithmic number of rounds.
@@
A key open question is understanding the distributed complexity landscape of the family of local potential problems. When $\Delta=O(1)$, our work shows that there are local potential problems with deterministic round complexity $\operatorname{log}^{\Theta(1)}n$, and by prior work there are also local potential problems with round complexity $\Theta(\operatorname{log}^{*}n)$ and $O(1)$. However, one question that is currently wide open is what happens when we step outside the deterministic LOCAL model. For example, does randomness ever help with any of these problems? The only lower bound we have is the exponentially lower $\Omega(\operatorname{log}\operatorname{log}n)$ for locally optimal cuts; is this the correct complexity? If yes, do all local potential problems admit $\operatorname{poly}(\operatorname{log}\operatorname{log}n)$ randomized algorithms? Do any of these problems admit a distributed quantum advantage?
Also, questions related to decidability are wide open. For example, given the specification of a local potential problem, is it decidable whether it can be solved in $O(\operatorname{log}^{*}n)$ deterministic rounds or whether it requires $\Omega(\operatorname{log}n)$ deterministic rounds? (We remind the reader of the well-known fact that, for LCL problems, there is a gap in the deterministic complexity landscape between $O(\operatorname{log}^{*}n)$ and $\Omega(\operatorname{log}n)$.)
@@ 1.3 New techniques and key ideas
Intuitively, the key challenge in our example was that the solution was very fragile—all nodes were barely happy, and a minor perturbation in their local neighborhood may cause them to flip, which leads to long cascades that we must avoid.
@@
Let us now try to identify a key combinatorial property of locally optimal cuts (and more generally, local potential problems), focusing on bounded-degree graphs for simplicity. Consider first a sequential algorithm that eliminates all unhappy nodes one by one, let $X$ be the set of all nodes that were flipped in this process at least once. Note that in general, $X$ might span the entire input graph.
@@
Our algorithm uses the MPX clustering algorithm [40, 28] to divide the nodes into polylogarithmic-radius clusters with small boundaries. Each cluster can make nodes far from boundaries safe; we can simply compute for each cluster the result of repeatedly finding for each internal node an improving set $A$, and flipping $A$ until all internal nodes are safe.
@@
Our algorithm is randomized, but it can be derandomized (with polylogarithmic cost) by using standard techniques [33, 44, 32].
Lower bound (Section 5).
If we parameterize the running time of our algorithm both as a function of maximum degree $\Delta$ and the number of nodes $n$, we have an algorithm for locally optimal cuts that runs in $O(\Delta^{2}\operatorname{log}^{6}n)$ rounds; in particular, the dependence on $n$ is only polylogarithmic, but the dependence on $\Delta$ is polynomial.
In Section 5 we prove, through an indistinguishability argument, that there is no (deterministic or randomized) algorithm for locally optimal cuts that solves the problem in time that is only polylogarithmic in both $n$ and $\Delta$; we prove a lower bound of $\Omega(\min\{\Delta,\sqrt{n}\})$ rounds.
For the high-level idea of the lower bound, observe that in the locally optimal cut problem, the output (1-1 or +1+1) of a node $v$ is determined if the outputs of all neighbors of $v$ are fixed and if the number of 1-1 and +1+1 outputs among the neighbors is not equal. In this case, $v$ must output the minority value among the neighbors. Hence, if $v$ has odd degree, the output of $v$ is determined by the outputs of the neighbors and thus, if two odd-degree nodes $v$ and $v^{\prime}$ have the same set of neighbors, then $v$ and $v^{\prime}$ must have the same output in any valid solution. Based on this observation, we build a path-like lower bound graph for which any valid locally optimal cut solution corresponds to a valid proper $2$-coloring of a path, a problem that is known to have a round complexity lower bound that is linear in the length of the path [39].
The construction consists of a sequence of layers such that each layer forms an independent set and consecutive layers are connected by a complete bipartite subgraph. Note that all nodes in the same layer then have the same set of neighbors. The layers are constructed such that all nodes have an odd degree and hence all nodes in the same layer must output the same value. Further, the layers are chosen in such a way that for any two consecutive layers, the outputs must differ. As a consequence, any locally optimal cut solution on the layered graph induces a $2$-coloring of the underlying path. Finally, we make sure that the construction is symmetric enough so that the $2$-coloring can only be produced if the endpoints of the path coordinate with each other. The claimed lower bound then follows because for a sequence of length $k$, we need $\Theta(k^{2})$ nodes and a maximum degree $\Delta$ of order $\Theta(k)$.
Since the indistinguishability argument has been proved to work also in the quantum-LOCAL model (and, actually, even the stronger non-signaling model [31, 2, 22]), the same lower bound applies to quantum algorithms as well.